OpenStudy (anonymous):

Help...I am trying to solve this problem. Quadratic Formula 9x2-18x+3=0 This is what I have so far. -(18)plus or minus 6 square 6 over 18. This is the way the instructor is teaching us. Can anyone help?

OpenStudy (anonymous):

-(18) \[-(18)\pm 6\sqrt{6} \over 18\]

OpenStudy (radar):

\[-3(+ \- )\sqrt{6}\]

OpenStudy (anonymous):

Thank you, but how did you get the answer. Did you divide -18 by 6?

OpenStudy (radar):

I divided all terms by 3 then a=3, b=-6, c=1, I messed up using it in the equation. I will try again: (b^2-4ac)^1/2 becomes (36-4x3x1)^1/2=(24)^1/2=(4x6)^1/2=2(6)^1/2 now plugging it in it becomes (6+/-2(6)^1/2)/6 so the answer would be

OpenStudy (radar):

\[(3=/-\sqrt{6})/3\]

OpenStudy (radar):

change that equal sign to a +

OpenStudy (radar):

plus or minus the square root of 6